Galvanized nails are iron nails that have been plated with zinc to prevent rusting. The relevant reaction is

Zn2+(aq)+2e−→Zn(s)

For a large batch of nails, a manufacturer needs to plate a total zinc mass of 3.20 kg on the surface to get adequate coverage.

Part B:

How many coulombs of charge are needed to produce 48.9 mol of solid zinc?

Express your answer numerically in coulombs.
Answer #1

1 mol of Zn needs 2 mol of electrons

So,

48.9 mol of Zn would need = 2*48.9 = 97.8 mol of electrons

Number of electrons = mol of electrons * Avogadro’s number

= 97.8 * 6.022*10^23

= 5.89*10^25 electrons

Charge required = Number of electrons * charge on one electron

= 5.89*10^25 * 1.602*10^-19 C

= 9.44*10^6 c

Answer: 9.44*10^6 c